### v1

# Token Pruning in Audio Transformers: Optimizing Performance and Decoding Patch Importance

Abstract
Vision Transformers (ViTs) have achieved state-of-the-art performance across various computer vision tasks, but their high computational cost remains a challenge. Token pruning has been proposed to reduce this cost by selectively removing less important tokens. While effective in vision tasks by discarding non-object regions, applying this technique to audio tasks presents unique challenges, as distinguishing relevant from irrelevant regions in time-frequency representations is less straightforward. In this study, for the first time, we applied token pruning to ViT-based audio classification models using Mel-spectrograms and analyzed the trade-offs between model performance and computational cost: TopK token pruning can reduce MAC operations of AudioMAE and AST by 30-40%, with less than a 1% drop in classification accuracy. Our analysis reveals that while high-intensity tokens contribute significantly to model accuracy, low-intensity tokens remain important. In particular, they play a more critical role in general audio classification tasks than in speech-specific tasks.
Index Terms:
Audio classification, Audio spectrogram transformers, Token pruning.I Introduction
The Vision Transformer (ViT) [1] has achieved various SOTA (state-of-the-art) results in many downstream tasks. As the Transformer [2] has data-agnostic characteristics and audio can be represented as 2D data using Mel-spectrograms, previous studies [3, 4, 5, 6, 7, 8, 9, 10] have shown the applicability of ViT to audio downstream tasks.
To reduce the high computational demands of Transformer-based models, token reduction methods have been proposed, as the number of tokens is a quadratic factor in both time and memory complexity. In image classification tasks, various token pruning methods based on attention scores [11, 12, 13] or DNN predictors [14, 15] have demonstrated favorable trade-offs between accuracy and computational cost. As shown in Fig. 1, these methods gradually remove tokens not related to the object or background patches throughout the pruning stages. This is a reasonable selection because such patches are not necessary for classifying the object to be classified.
However, in audio classification tasks, it is unclear which tokens should be discarded: empty or low-intensity regions in Mel-spectrograms do not necessarily indicate a lack of information. In the task of identifying the sound of a baby crying, losing the brief silences between cries could make it more difficult to distinguish it from the sound of a siren. Furthermore, empty regions in certain frequency bands of a Mel-spectrogram can be a characteristic of the sound source, which should not be ignored.
The works most similar to ours are [16] and [17], which apply token merging for general audio classification and TopK pruning for speech information retrieval with speech LLM, respectively. Unlike previous research, our study focuses on analyzing which kind of tokens are important for prediction in ViT-based audio classification models when token pruning is applied. Our main findings are summarized as follows:
TopK token pruning based on attention scores can reduce the Multiply-Accumulate Count (MAC) operations of AudioMAE and AST by 30-40%, with less than a 1% drop in accuracy.
We visualize pruning patterns in audio models and observe that they effectively discard pauses and padding, while selectively retaining tokens from low-intensity or low-complexity Mel-spectrogram patches.
We measure Kendall’s between attention scores and statistical features (e.g., the intensity (mean) and variation (std) of signals in the patches) and found strong alignment. While pruning based on these statistics performs comparably, attention-based pruning achieves better results, suggesting models benefit from preserving less strong and complex acoustic regions.
Using selective token pruning based on intensity, we show that while high-intensity tokens contribute significantly to model performance, low-intensity tokens are more important for general audio classification tasks (e.g., AS-20K, ESC-50) than for speech-specific tasks (e.g., SPC-2, VoxCeleb-1).
- baseline, : keep-rate, - attention score, : intensity (), : variation ().
| - | AS-20K | SPC-2 | ESC-50 | VoxCeleb-1 | |||||||||||||||||
| AudioMAE | AST | AudioMAE | AST | AudioMAE | AST | AudioMAE | |||||||||||||||
| 37.8 | 38.7 | 98.36 | 97.33 | 94.10 | 95.05 | 95.18 | |||||||||||||||
| kr | |||||||||||||||||||||
| 0.9 | 37.3 | 36.7 | 37.0 | 38.7 | 36.6 | 36.8 | 98.09 | 98.27 | 98.33 | 97.21 | 97.08 | 97.06 | 93.66 | 93.39 | 93.75 | 94.36 | 92.05 | 92.60 | 95.05 | 95.16 | 95.21 |
| 0.8 | 36.8 | 35.7 | 36.5 | 37.9 | 35.9 | 36.2 | 97.77 | 98.26 | 98.38 | 97.22 | 97.11 | 97.14 | 93.45 | 92.97 | 93.67 | 94.32 | 91.80 | 92.32 | 94.77 | 94.66 | 95.27 |
| 0.7 | 36.2 | 34.5 | 35.5 | 37.8 | 35.1 | 35.7 | 97.70 | 98.18 | 98.24 | 97.19 | 96.77 | 96.98 | 93.46 | 92.25 | 93.16 | 94.37 | 91.03 | 92.13 | 94.46 | 93.37 | 94.64 |
| 0.6 | 35.5 | 33.0 | 34.1 | 37.6 | 34.4 | 34.9 | 97.66 | 97.95 | 98.20 | 97.20 | 96.82 | 96.81 | 93.16 | 91.79 | 92.40 | 94.37 | 90.31 | 91.48 | 93.38 | 90.47 | 93.06 |
| 0.5 | 34.4 | 31.0 | 32.4 | 37.2 | 33.4 | 34.0 | 97.44 | 97.68 | 97.85 | 97.11 | 96.72 | 96.74 | 92.75 | 90.79 | 91.23 | 94.07 | 89.36 | 89.54 | 91.26 | 86.58 | 89.54 |
| 0.4 | 32.8 | 28.4 | 30.2 | 36.8 | 32.1 | 32.7 | 97.35 | 97.48 | 97.35 | 97.07 | 96.60 | 96.84 | 91.87 | 89.53 | 90.02 | 93.87 | 87.69 | 88.37 | 88.02 | 80.81 | 84.71 |
II Token Pruning on Audio Transformer Models
II-A TopK Token Pruning on Audio Classification Transformers
To apply token pruning, we adopted AudioMAE and AST as they are representative audio classification models trained by masked auto-encoding and supervised training, respectively. We use TopK as a token pruning method since it is a competitive method and allows us to distinguish the origin of tokens [18]. After the raw waveform is converted into a Mel-spectrogram, it is treated as an image so that patch embeddings and positional embeddings can be applied. The token pruning module is placed between the multi-head self-attention module and the MLP module of selected (the 4th, 7th, and 10th) ViT blocks as in Fig. 2. The choice of pruning location follows the previous token pruning works [11, 12, 13, 14]. In AudioMAE, we used token-to-token attention scores in (2a), which quantify the attention each token receives from others as shown in (1) - as indicators of token importance since mean pooling is used for the final prediction. In AST, we used CLS attention scores (2b), following [11]. Among the tokens from Mel-spectrogram, we retained tokens with the highest scores in each pruning block. The same keep-rate is applied to all pruning-enabled blocks.
II-B Datasets and Metrics
We evaluated audio models using AudioSet Balanced (AS-20K) [19], Speech Commands V2 (SPC-2) [20], Environmental Sound Classification (ESC-50) [21] and VoxCeleb-1 [22]. The number of classes is 527, 35, 50 and 1251 respectively. We reported the maximum mAP for AS-20K and Top-1 accuracy for other datasets. Since AST shows low accuracy on VoxCeleb-1 (30.1) [4], we excluded it from our experiments.
| Dataset | 1.0 | 0.9 | 0.8 | 0.7 | 0.6 | 0.5 | 0.4 | |
| SPC-2 | 64 | 5.6 | 4.9 | 4.30 | 3.7 | 3.3 | 2.8 | 2.5 |
| ESC-50 | 256 | 23.1 | 20.0 | 17.3 | 15.0 | 13.1 | 11.4 | 10.0 |
| AS-20K VoxCeleb-1 | 512 | 48.6 | 41.8 | 36.0 | 31.1 | 27.1 | 23.7 | 20.8 |
II-C Training Hyperparameters
We downloaded checkpoints of AudioMAE and AST pre-trained on AudioSet-2M, using a ViT-B configuration with -sized patches without strides. For keep-rate scheduling we adopted the method used in EViT [11]. We followed the original training procedures and outline the modified hyperparameters in the order of AS-20K, SPC-2, ESC-50, VoxCeleb-1. For AudioMAE, batch sizes are 16, 256, 64, 32 and the minimum learning rate is set to for all datasets. We trained AudioMAE for 60, 90, 120, 90 epochs for four benchmarks, with keep-rate reduction starting at the 30th, 10th, 20th, 20th epoch and continuing for 20, 30, 40, 40 epochs. Before token pruning starts, we applied masking ratio 0.3 for AS-20K and ESC-50; 0.0 for others. Once pruning is enabled, all masking strategies including SpecAug [23] were disabled, since pruning itself already applies strong masking. For AST, batch sizes were set to 64, 128, 48 and learning rates to , , . We trained AST for 30 epochs for all dataset, enabling pruning at the 15th, 5th, 5th epoch and reducing keep-rate for 10, 15, 15 epochs. We used PyTorch [24] with two GPUs and automatic mixed precision [25] for our training.
II-D Benchmark Results of Token Pruning on Audio Models
Table I and II demonstrate that a simple TopK pruning based on attention scores can reduce the computation (MAC) by 30-40%, with less than a 1% drop in accuracy with both audio transformer models. We also found that AudioMAE is more sensitive to token loss than AST especially at lower keep‑rates, likely due to its reliance on multiple tokens for the final prediction. Furthermore, accuracy tends to drop more when more tokens are pruned—particularly for harder tasks with a larger number of classes.
III Analysis of TopK Pruning on Audio Models
III-A Visualization of Pruning Patterns on Audio Models
We begin our analysis of important tokens by visualizing TopK token pruning patterns of AudioMAE and AST in Fig. 3. The models effectively discards tokens originating not only from the padding regions (A) but also from pauses between sounds (B). This observation suggests that models preferentially retain tokens from high-intensity Mel-spectrogram patches. However, some tokens from low-intensity regions are retained even if they are not part of padding or pause regions. Additionally, some high-intensity tokens are discarded instead of low-intensity ones (A, C, D), indicating that they received lower attention scores than certain low-intensity patches.
III-B Token Pruning with Statistical Features
Based on this observation, we wondered whether we can prune tokens using features directly available from the Mel-spectrogram. Specifically, we considered the intensity ( of signals in a patch), as high-intensity patches typically carry more energy, and variation (), as it can capture the patch’s texture complexity in the Mel-spectrogram. This motivated us to explore whether these statistics could serve as alternative pruning metrics to attention scores. We fine-tuned both models using and as pruning metrics, replacing attention scores while keeping the same hyperparameters. In Table I, these statistics perform as well as attention scores at high keep-rates (e.g., 0.7), but fail to outperform attention scores when tokens are aggressively pruned. This suggests that audio models also needs to retain low-intensity or little-variation patches for better prediction.
| Dataset | C1 | C2 | C3 | C4 | C5 |
| AS-20K | 23.1 | 43.2 | 20.4 | 11.9 | 1.5 |
| SPC-2 | 18.5 | 33.8 | 27.5 | 15.0 | 5.2 |
| ESC-50 | 12.1 | 47.5 | 23.1 | 13.4 | 3.9 |
| VoxCeleb-1 | 32.9 | 23.5 | 24.7 | 14.1 | 4.8 |
III-C Correlation Between Attention and Statistical Features
As shown in [13], in image classification tasks ViT performs better if tokens corresponding to objects are not pruned. This is reasonable because clear boundaries exist between objects and the background in images. However, it is hard to say that audio models prunes tokens well merely by observing the pattern since low-intensity regions might still contribute to class discrimination. To investigate the models’ preference over high-intensity regions, we quantify the preference using Kendall’s correlation [26] between intensity and variation of each patch and the corresponding attention score. Clustering is necessary to avoid ranking reordering due to minor differences in intensity, which is not perceptually differentiable. Fig. 4 shows that both statistics are positively correlated with attention score, which explains the comparable performance to the attention based TopK model.
| block | AudioMAE | AST | |||||
| AS-20K | SPC-2 | ESC-50 | Vox-1 | AS-20K | SPC-2 | ESC-50 | |
| 35.9 | 97.43 | 91.54 | 94.39 | 35.5 | 95.36 | 88.56 | |
| 35.8 | 97.62 | 91.24 | 94.62 | 36.6 | 96.56 | 92.49 | |
| 36.0 | 97.75 | 91.86 | 94.68 | 37.9 | 97.22 | 94.69 | |
| 36.4 | 97.91 | 92.26 | 94.91 | 38.4 | 97.31 | 94.83 | |
| 36.5 | 98.10 | 92.75 | 95.21 | 38.6 | 97.35 | 94.93 | |
| 37.2 | 98.25 | 93.05 | 95.62 | 38.7 | 97.32 | 94.98 | |
| 27.4 | 71.14 | 71.71 | 66.17 | 19.5 | 44.74 | 54.50 | |
| 27.1 | 71.30 | 72.55 | 66.88 | 21.7 | 63.50 | 57.13 | |
| 26.8 | 74.50 | 72.94 | 67.66 | 24.9 | 94.40 | 67.66 | |
| 28.0 | 78.54 | 75.78 | 73.36 | 27.5 | 97.10 | 79.65 | |
| 30.1 | 88.84 | 81.19 | 86.37 | 28.9 | 97.30 | 82.24 | |
| 35.3 | 98.22 | 91.52 | 94.13 | 33.2 | 97.30 | 90.93 | |
III-D Impact of removing tokens from low/high-intensity regions
We assess the impact of discarding tokens belonging to low- (C1, C2: ) or high-intensity (C4, C5: ) groups on accuracy during inference. In this test, the tokens belonging to these groups are pruned after being processed by specific blocks. In Table IV, and indicate the block index at which low- and high-intensity token groups are removed, respectively. High-intensity tokens contribute significantly to model performance. In addition, low-intensity tokens become more important for general audio classification tasks than for speech tasks. AudioMAE shows greater robustness to the loss of both token types across all datasets compared to AST.
III-E Inspection of retained patches’ statistics
We visualize the relationship between the and of signals in patches for each dataset in Fig. 6. The vertical lines on low-mean patches in the input Mel-spectrograms of ESC-50 and SPC-2 indicate artifacts from padding regions at the end of the audio samples; AudioMAE effectively discards these regions, whereas AST does not. Except for SPC-2, which consists of a single word, there are two patch groups on either side of the diagonal in each histogram. The retention of these groups after pruning indicates the audio model’s reliance on (1) low-to-mid intensity patches with greater complexity and (2) high-intensity patches with less complexity. For general audio classification tasks (AS-20K, ESC-50), we observe that AudioMAE retains significantly more low-intensity tokens compared to AST. AudioMAE retains approximately 1.9× more tokens belonging to low-intensity clusters (C1 and C2) after pruning than AST in both datasets. While both models and datasets retain empty (bottom-left most) patches after pruning, AudioMAE prunes all of those patches in the speaker identification task (VoxCeleb-1).
IV Conclusion
In this work, we show that TopK token pruning can be effectively applied to audio transformer models in classification tasks, achieving a competitive trade-off between accuracy and computational cost. We explore using intensity and variation of signals in patches as alternative token importance indicators, supported by positive Kendall correlations with attention scores. However, attention-based pruning consistently outperforms these statistics, suggesting that both low-intensity and low-variation patches are important. Visualization and ablation studies confirm that the model attends to such patches, especially in general audio classification tasks, highlighting their non-negligible contribution to higher classification accuracy.